OpenStudy (anonymous):

A farmer wants to build a fence to enclose an area 12 feet long and 7 feet wide. How many feet of fencing will be needed? [A] 31 feet [B] 168 feet [C] 38 feet [D] 84 feet

OpenStudy (anonymous):

P= 2(L+W) P= 2(7+12) P= 38